OpenStudy (anonymous):

Which of the following is greatest? Choose one answer. a. eight twenty-fifths b. zero point three three c. forty-one one hundredths d. seventeen-fiftieths

OpenStudy (anonymous):

c

OpenStudy (anonymous):

ty :)

OpenStudy (anonymous):

\[\frac{ 8 }{ 25 }\times \frac{ 4 }{ 4 }=\frac{ 32 }{ 100 }=0.32\] 0.33 \[\frac{ 41 }{ 100 }=0.41\] \[\frac{ 17 }{ 50 }\times \frac{ 2 }{ 2 }=\frac{ 34 }{ 100 }=0.34\] you can judge ,which is greater.

OpenStudy (anonymous):

so c is correct?